OpenStudy (anonymous):

Find the unit vector that has the same direction as the vector v. v=6i-8j

OpenStudy (anonymous):

divide by the length of the vector

OpenStudy (anonymous):

which is \[|+i-8j|=\sqrt{6^3+8^2}\]

OpenStudy (anonymous):

Why is it 6 to the third power?

OpenStudy (anonymous):

because i made a typo it is \[\sqrt{6^2+8^2}=10\]

OpenStudy (anonymous):

divide by 10 and get \[.6i-.8j\]

OpenStudy (anonymous):

Got it! thank you!

OpenStudy (anonymous):

Write the vector v in terms of i and j whose magnitude and direction angle are given. lvl=6/7, theta =118degrees The answer I got is 0.24i+0.75j is that right?

OpenStudy (anonymous):

*-0.24